paul left springfield traveling north at an average speed of 64 miles per hour. tim left springfield traveling south at an avergae spped of 59 miles per hour. how long would it take after paul left for them to be 743 miles apart? how many hours would it be after paul left?
Answer by stanbon(75887)

paul left springfield traveling north at an average speed of 64 miles per hour. tim left springfield traveling south at an avergae speed of 59 miles per hour. how long would it take after paul left for them to be 743 miles apart? how many hours would it be after paul left?
---------------------------
They are separating at a rate of (64+59) = 123 mph
---
time = distance/rate = 743/123 = 6.04 hrs (same time for both.
---
Note: 0.04 hr = 0.04*60 minutes is approximately 3 minutes
---
Ans: Approximate time = 6 hr. 3 minutes
============================================
Cheers,
Stan H.